# Multimodal Fusion on Low-quality Data: A Comprehensive Survey

### v1

@@ 1 Introduction
Our perception of the world is based on multiple modalities, e.g., touch, sight, hearing, smell and taste. Even when some sensory signals are unreliable, humans can extract useful clues from imperfect multimodal inputs and further piece together the whole scene of happening events [1]. With the development of sensory technology, we can easily collect diverse forms of data for analysis. To fully release the value of each modality, multimodal fusion emerges as a promising paradigm to obtain precise and reliable predictions by integrating all available cues for downstream analysis tasks, e.g., medical image analysis, autonomous vehicles [2, 3] and sentiment recognition [4, 5, 6]. Intuitively, fusing information from different modalities offers the possibility of exploring cross-modal correlation and gaining better performance. However, there is growing recognition that widely-used AI models are often mislead by spurious correlations and biases within the low-quality data. In real-world setting, the quality of different modalities usually varies due to unexpected environmental factors or sensor issues. Some recent studies have shown both empirically and theoretically that conventional multimodal fusion may fail on low-quality multimodal data in the wild, e.g., imbalanced [7, 8, 9, 10], noisy [11] or even corrupted [12] multimodal data. To overcome such a limitation and take a step to strong and generalized multimodal learning in the real world, we identify the properties of low-quality multimodal data and focus on some unique challenges for multimodal machine fusion in real-world setting. We also highlight technical advances that may help to make multimodal fusion more reliable and trustworthy in the open environment. In this paper we identify and explore four core technical challenges surrounding multimodal fusion on low-quality multimodal data. They are summarized as follows (also intuitively illustrated in Fig. 1):
@@
    Incomplete multimodal data. The second fundamental challenge is learning with incomplete multimodal data (i.e., multimodal data with partial missing modalities). For instance, in the medical field, patients even with the same disease may choose different medical examinations producing incomplete multimodal data. Developing flexible and reliable multimodal learning methods that can handle incomplete multimodal data is a challenging yet promising research direction.
@@
    Quality dynamically varying multimodal data. The forth fundamental challenge is how to adapt the quality dynamically varying nature of multimodal data. In practice, the quality of one modality often varies for different samples due to unforeseeable environment factors or sensor issues. For example, the RGB image is less informative than the thermal modality in low-light or backlight conditions. Therefore, dynamically integrating multimodal data by being aware of the varying quality in fusion is necessary in real-world applications.
Towards addressing these increasingly important issues of multimodal fusion, this study systematically organizes the key challenges through several taxonomies. Different from previous related works discuss about various multimodal learning tasks [13, 14], this survey is mainly focusing on multimodal fusion which is the most fundamental problem in multimodal learning and the unique challenges raising upon low-quality multimodal data in downstream tasks, including clustering, classification, object detection and semantic segmentation. In the following sections, we detailed this field by recent advances and technical challenges of multimodal fusion facing low quality data: learning on noisy multimodal data (Section 2), missing modality imputation (Section 3), balanced multimodal fusion (Section 4), and dynamic multimodal fusion (Section 5). A discussion in Section 6 is provided as a conclusion.
@@ 2 Learning on Noisy Multimodal Data
Multimodal noise can be be roughly split into two categories according to their source: 1) modal-specific noise that arise from sensor errors, environmental factors, or transmission of each modality separately, 2) cross-modal noise that arise from unaligned multimodal pairs which can be regarded as semantic level noise.
@@ 2.1 Modal-specific noise reduction
Modal-specific noise reduction is highly relying on the input modalities and task at hand. In this section, we focus on visual noise reduction in multimodal image fusion task to present a showcase. Most modal-specific denoise methods focus on aggregating useful information from multimodal data and mitigate the influence of noise in multimodal fusion.
@@ 2.1.1 Weighted average fusion
A simple multimodal noise reduction way is to conduct average fusion over multimodal data. Due to the randomness of noise, the average operation effectively reduces the ratio of noise in fusion result. Rajalingam et al. [19] developed an intelligent multimodal fusion algorithm. The method decomposes the input image into high-frequency and low-frequency components and applied an average fusion rule for the low-frequency components while utilizing guided filtering for the high-frequency components. Essentially, average-based multimodal noise reduction methods assign the same fusion weights to different modalities. However, the noise severity of each modality is different and varies among multimodal samples. To address aforementioned issue, weighted average-based multimodal noise reduction method is proposed. Budhiraja et al. [20] proposed a multimodal image fusion scheme based on the guided filter. Their approach involves Gaussian decomposition with local average energy and average gradient-based salient maps for the base layer and detail layer, respectively. Qu et al. [21] proposed a fusion rule that used for calculating the wavelet transformation modulus maxima of input images with different bandwidths and levels. The approach aims to effectively combine the information from multiple image modalities. Achim et al.[22] developed a fusion rule based on fractional lower order moments for magnetic resonance (MR) and Computed Tomography (CT) images in the multi-scale medical image analysis literature. Considering the similarities and disparities between the different modalities, Yu et al. [23] divided the image into the common component and innovation component and represented the source image with the common and innovation sparse coefficients by joint sparse representation. Wang et al. [24] designed a gate mechanism to dynamically fuse visual features. By utilizing visual features as supplementary information, it is able to effectively mitigate the influence of noise in voice signals. However, this strategy assumes that the visual data is clean. If the visual data is also corrupted by noise, this solution may be sub-optimal. Sadeghi et al. [25] adopted a latent sequential variable with Markovian dependencies to decide whether visual data is beneficial to audio-visual speech enhancement. Bai et al.[18] proposed a soft-association mechanism to handle inferior image conditions. In practice, the self-attention in Transformer enables model to adaptively decide where and what information should be taken from the image. Cheng et al. [16] adopted
both soft attention and hard attention to mitigate noise issue, which is caused by rolling-shutter effect in lidar-stereo fusion task. Specifically, a fusion block is utilized to aggregate supplement information from stereo cameras and generates two disparity maps. Then, through comparing disparity map and lidar point, the noisy points in raw data are filtered out.
@@ 2.1.2 Joint variation based fusion
Joint optimization is often used in variation based noise reduction when fusing multiple visual modalities, e.g., RGB and thermal. In unimodal noise reduction field, taking image modality as an example, classical denoise model based on Total Variation (TV) is equivalent to solving the following optimization problems:
@@
where $u_{m}(\boldsymbol{x})$ represents the input data from modality $m$, $w_{m}$ is the non-negative weight satisfying $\sum_{m=1}^{M}w_{m}=1$. In a nutshell, joint optimization is the solution of a specific optimization problem using data from different modalities. Wang et al. [26] proposed a variational model in the pixel domain and wavelet domain for fusion and denoise of noisy multifocus images. Kumar et al. [27] also employed the total variational model to fuse images acquired using multiple sensors and achieves excellent results on two modal images provided in the medical field and aircraft navigation field. Recently, Padmavathi et al. [28] proposed an image fusion algorithm for constructing a fused image using the total variation model. They incorporated an optimized adaptive weighting fusion scheme for medical multimodal data analysis (i.e., Magnetic Resonance images and Positron Emission Tomography images). Nie et al. [29] introduced a total variation-based fusion method for infrared and visible images. The method leverages the total variation model to achieve fusion of the two modalities, enabling enhanced image quality and information integration. Quan et al. [30] developed a feature fusion method called relative total variation structure analysis (RTVSA) for urban area classification. This approach combines various features derived from hyperspectral imaging (HSI) and LiDAR data, providing a comprehensive analysis for accurate classification. Furthermore, Liu et al. [31] designed a two-stage enhancement (TSE) framework for infrared and visible image fusion. Their framework incorporated an attention mechanism and a feature-linking model (FLM), combining structure adaptive total-variational (SATV) and L1 sparsity terms to extract two-scale detail layers and a base layer. The proposed method demonstrates robustness and effectiveness in improving the quality of fused images.
@@ 2.2 Cross-modal noise reduction
Many multimodal tasks (e.g., multimodal object detection, vision-and-language understanding) are highly relying on correctly aligned multimodal training data. However, the large-scale web-crawled multimodal pairs tend to contain weakly aligned or even unaligned samples [32]. For example, in multimodal object detection, image pairs are commonly suffer from weakly aligned issue, i.e., the locations of the same objects may be shifted across different modalities [33]. In social media, the text descriptions are frequently irrelevant with the image content, i.e., unpaired modalities. In this section, we consider weakly aligned or unaligned multimodal samples as cross-modal noise. Compared to modal-specific noise, cross-modal noise lies in a higher-level semantic space. Current cross-modal noise reduction methods can be roughly split into rules-based filtering, model-based rectifying and noise robustness regularization.
@@
Noise robust regularization is another perspective to mitigate the influence of cross-modal noise. To explicitly stabilize and harmonize multimodal pretraining and mitigate the influence of latent noise, NLIP [41] employs noise-adaptive regularization to avoid overfitting to noisy image-text pairs, adjusting alignment labels based on estimated noisy probabilities. Additionally, NLIP utilizes a concept-conditioned cross-modal decoder to generate synthetic captions to impute the missing object information. To mitigate the influence of cross-modal noise in vision-language task, Li et al. [42] proposed OSCAR that detects object tags in images and uses them as anchor points for alignment. Recent work provides a theoretical and empirical to shed light on the influence of cross-modal noise in multimodal contrastive learning [43]. Furthermore, the research introduces a novel MMCL loss designed to handle unpaired multimodal samples, demonstrating enhanced robustness in empirical experiments.
@@ 3 Incomplete multimodal Learning
The collected multimodal data in real applications are often incomplete with partial missing modalities for some samples due to unexpected factors, such as equipment damage, data transmission and storage loss. For a concrete example, in user-oriented recommendation system, the browsing behavior history and credit score information may not always be available for some users [48]. Similarly, although combining data from multiple modalities e.g., magnetic resonance imaging (MRI) scans, positron emission tomography (PET) scans, and cerebrospinal fluid (CSF) information, can obtain a more accurate diagnosis for Alzheimer’s Disease [49, 50], owing to high measurement cost for PET scans and uncomfortable invasive tests for CSF, some patients may refuse to conduct these examinations. Thus, incomplete multimodal data commonly exist in Alzheimer’s Disease diagnosis [51]. Typically, conventional multimodal learning models assume the completeness of multimodal data and thus cannot be applicable to the case with partial missing modalities directly. For this issue, incomplete multimodal learning, aiming at exploring information of incomplete multimodal data with partial missing modalities, appears and has received increasing research attentions in recent years [52]. In this section, we mainly focus on the current progress of researches on incomplete multimodal learning. From the view of whether to impute the missing data, we categorize existing methods into two groups including imputation based and imputation-free based incomplete multimodal learning, where the imputation based methods are further divided into two groups as shown in Fig. 2 including instance and modality level imputation.
@@ 3.1 Imputation based incomplete multimodal learning
The power of multimodal learning has been well recognized on the strictly aligned multiple modalities. For the incomplete multimodal data, a natural approach to address the missing modality learning issue is to directly impute the missing modalities as shown in Fig. 2(c) or related missing information in the kernels or graphs constructed from the data as shown in Fig. 2(b). Then the conventional multimodal learning algorithm can be applied to downstream tasks. In this branch, the existing imputation based works can be further divided into two categories, including model-agnostic imputation and model-specific imputation. The model-specific imputation methods generally design a imputation model or network for missing data recovery while the imputation model is not necessary for the model-agnostic imputation methods which usually utilize heuristic filling strategy for the missing modalities.
@@ 3.1.1 Model-agnostic imputation
For the missing modalities, the simplest imputation schemes are zero imputation and mean value imputation, i.e., filling missing values with zero or imputation with the averaged value in the corresponding modality, which are widely used as the baseline methods in most studies [53, 54]. However, previous work [53] shows that the naive imputation may lead to unpromising performance. To address this issue, a weighted matrix factorization model is proposed to assign small weights for the imputed modalities, mitigating the negative influence on optimization [53]. Furthermore, for an online multimodal clustering case, a dynamic weighting and imputation mechanism is designed according to the valid modalities and task at hand[55]. To implement the graph based approach on incomplete multimodal data, Wang et al. [56] utilize similarity values computed from valid modalities to impute the missing elements of graph. In [57], the missing elements in the incomplete kernels are pre-imputed by the average of all columns. In [58], the authors also utilize the average instance of the corresponding modality to impute the missing instances and then assign a small weight to those imputed instances during the graph construction phase.
@@ 3.1.2 Learning based imputation
For the learning based imputation, according to the filling values, we can further divide it into two categories: kernel/graph based imputation and raw data imputation.
Kernel/Graph based imputation: Kernel and graph learning based approaches are two main lines of incomplete multimodal learning, which transform the learning phase from the raw data to the kernel space or graph. For the kernel and graph constructed from each modality, each element represents the relationship in kernel space or affinity/similarity of instance pairs. The partial missing instances in each modality result in the incomplete kernels and graphs. A line of works have been proposed to address the incomplete kernel/graph based multimodal learning issue. Here we highlight a few seminal works. David Williams et al. [59] proposed a Gaussian mixture model based method that can analytically compute the missing Gaussian kernel elements associated with the missing instances. Trivedi et al. [60] proposed a kernelized canonical correlation analysis (KCCA) based imputation method, which minimizes a Laplacian kernel regularization to obtain the missing kernel elements. Although effective, it is only applicable to the two-modal data with one complete modality. Lately, collective kernel learning (CKL) [48] is proposed to mutually complete the two incomplete kernels associated with the two incomplete modalities. For the data with more than two modalities, such as the case with $k$ modalities ($k\geq 3$), CKL needs to recover each kernel matrix ${K_{i+1}}$ of the $\left({i+1}\right)$th modality according to kernel ${K_{i}}$ in a cycle iteration scheme. Sahely et al. [61] proposed a multimodal kernel completion method based on the sparse reconstruction technique. The method can complete the missing rows and columns of all kernels jointly in a within-kernel preservation and between-kernel exploration framework. Yang et al. reduce incomplete multi-modal to semi-supervised learning problem, and further propose SLIM (Semi-supervised Learning with Incomplete Modality ) and its kernel vision SLIM-K based on matrix completion for solving the modal incompleteness [62]. Recently, Liu et al. proposed several flexible kernel methods to recover the missing kernel elements for incomplete multimodal clustering tasks [63, 64, 65, 66]. For instance, MKKM-IK [63] unifies the kernel completion and consensus partition matrix learning in a unified framework:
@@
By aligning all incomplete kernels to a consensus kernel derived from the consensus orthogonal representation, these methods can jointly obtain the missing elements associated with missing instances and the consensus partition representation shared by all modalities for incomplete multimodal data clustering. Besides of kernel based methods, an adaptive graph completion method is developed to recover the missing similarities of all graphs with respect to all modalities [67]. The method mainly employs the sparse-representation-based data reconstruction technique to borrow the graph information of other modalities in recovering the missing elements of each modality.
@@
Raw data imputation: Some methods recover the missing modalities at raw feature level. For example, VIGAN (View Imputation via Generative Adversarial Network) [68] is a pioneering work based on generative adversarial network (GAN) and Autoencoder for missing modality recovery. VIGAN employs GAN to initialize the missing modality according to another modality of the same sample and then uses denoise Autoencoder to refine the recovered missing modalities. A limitation of VIGAN is that it is not applicable to data with more than two modalities. Subsequently, partial multimodal clustering via consistent GAN (PMVC_CGAN) [69] and adversarial incomplete multimodal clustering (AIMC) [70] are proposed for missing modality imputation and incomplete multimodal clustering. Different from VIGAN which uses one available data to generate the corresponding missing modality, PMVC_CGAN and AIMC seek to learn the missing modalities from the common representation encoded by the un-missing modalities of the same sample. In [71], a two-stage incomplete deep multimodal network is proposed for breast cancer prediction. Similar to the network structures of AIMC and PMVC_CGAN, GAN is introduced to enhance the quality of missing view imputation. Besides of GAN based methods, CRA (cascaded residual autoencoder) stacks all modalities into one modal and treats the missing modalities imputation as conventional missing feature completion by optimizing a single-modal cascaded residual autoencoder network [72]. Liu et al. [73] proposed another representative autoencoer based imputation framework termed RecFormer for clustering task. Specifically, RecFormer employs a two-stage autoencoder network with the self-attention structure to synchronously extract high-level semantic representations and recover the missing data
@@
To obtain a more reliable modality imputation performance, Tang et al. [74] proposed to impute the modality as the average of corresponding semantic neighbors. Lin et al. [75] established a dual prediction model that can predict the latent representation of the missing modality according to that of the observed modality and then used the decoder to obtain the recovered missing modality according to the predicted latent representation. matrix factorization is also applied for raw modality imputation. For example, UEAF (Unified Embedding Alignment Framework) [76] and IMCRV (Incomplete multimodal Clustering with Reconstructed Views) [77] provide two matrix factorization based missing modality recovery models, which can reversely reconstruct the data of missing modalities from the common representation. In [78], low-rank representation and tensor learning are introduced for joint missing modality recovery and graph completion. Explicitly taking the reliability of imputation into account, Xie et al. [79] introduces the Uncertainty-induced Incomplete Multi-View Data Classification (UIMC) model to address challenges in classifying incomplete multi-view data with missing perspectives. UIMC explores and exploits uncertainty by imputing missing modalities as distributions rather than deterministic values, allowing for more reliable predictions. The model employs an evidential multimodal learning framework to weigh imputed data, reducing the impact of low-quality imputations.

### v2

@@ 1 Introduction
Our perception of the world is based on multiple modalities, e.g., touch, sight, hearing, smell and taste. Even when some sensory signals are unreliable, humans can extract useful clues from imperfect multimodal inputs and further piece together the whole scene of happening events [1]. With the development of sensory technology, we can easily collect diverse forms of data for analysis. To fully release the value of each modality, multimodal fusion emerges as a promising paradigm to obtain precise and reliable predictions by integrating all available cues for downstream analysis tasks, e.g., medical image analysis, autonomous vehicles [2, 3] and sentiment recognition [4, 5, 6]. Intuitively, fusing information from different modalities offers the possibility of exploring cross-modal correlation and gaining better performance. However, there is growing recognition that widely-used AI models are often misled by spurious correlations and biases within low-quality data. In real-world setting, the quality of different modalities usually varies due to unexpected environmental factors or sensor issues. Some recent studies have shown both empirically and theoretically that conventional multimodal fusion may fail on low-quality multimodal data in the wild, e.g., imbalanced [7, 8, 9, 10], noisy [11] or even corrupted [12] multimodal data. To overcome such a limitation and take a step to strong and generalized multimodal learning in the real world, we identify the properties of low-quality multimodal data and focus on some unique challenges for multimodal machine fusion in real-world setting. We also highlight technical advances that may help to make multimodal fusion more reliable and trustworthy in the open environment. In this paper, we identify and explore four core technical challenges surrounding multimodal fusion on low-quality multimodal data. They are summarized as follows (also intuitively illustrated in Fig. 1):
@@
    Incomplete multimodal data. The second fundamental challenge is learning with incomplete multimodal data. For instance, in the medical field, patients even with the same disease may choose different medical examinations producing incomplete multimodal data. Developing flexible and reliable multimodal learning methods that can handle incomplete multimodal data is a challenging yet promising research direction.
@@
    Quality dynamically varying multimodal data. The fourth fundamental challenge is how to adapt the quality dynamically varying nature of multimodal data. In practice, the quality of one modality often varies for different samples due to unforeseeable environmental factors or sensor issues. For example, the RGB image is less informative than the thermal modality in low-light or backlight conditions. Therefore, dynamically integrating multimodal data by being aware of the varying quality is necessary in real-world applications.
Towards addressing these increasingly important issues of multimodal fusion, this study systematically organizes the key challenges through several taxonomies. Different from previous related works discuss about various multimodal learning tasks [13, 14], this survey is mainly focusing on multimodal fusion which is the most fundamental problem in multimodal learning and the unique challenges raising upon low-quality multimodal data in downstream tasks, including clustering, classification, object detection and semantic segmentation. In the following sections, we detailed this field by recent advances and technical challenges of multimodal fusion facing low-quality data: learning on noisy multimodal data (Section 2), missing modality imputation (Section 3), balanced multimodal fusion (Section 4), and dynamic multimodal fusion (Section 5). A discussion in Section 6 is provided as a conclusion.
@@ 2 Learning on Noisy Multimodal Data
Multimodal noise can be roughly split into two categories according to their source: 1) modal-specific noise that arises from sensor errors, environmental factors, or transmission of each modality separately, 2) cross-modal noise that arises from weakly aligned or unaligned multimodal pairs which can be regarded as semantic level noise.
@@ 2.1 Modal-specific noise reduction
Modal-specific noise reduction methods highly depends on the input modalities and task at hand. In this section, we focus on visual noise reduction in multimodal image fusion task to present a showcase. Most modal-specific denoise methods focus on aggregating useful information from multimodal data and mitigating the influence of noise in multimodal fusion.
@@ 2.1.1 Weighted average fusion
A simple multimodal noise reduction way is to conduct average fusion over multimodal data. Due to the randomness of noise, the average operation effectively reduces the ratio of noise in fused images. Rajalingam et al. [19] developed an intelligent multimodal fusion algorithm. The method decomposes the input image into high-frequency and low-frequency components and proposes an average fusion rule for the low-frequency components while utilizing guided filtering for the high-frequency components. Essentially, average-based multimodal noise reduction methods assign the same fusion weights to different modalities. However, the noise severity of each modality is different and varies among multimodal samples. To address the aforementioned issue, weighted average-based multimodal noise reduction method is proposed. Budhiraja et al. [20] proposed a multimodal image fusion scheme based on the guided filter. Their approach involves Gaussian decomposition with local average energy and average gradient-based salient maps for the base layer and detail layer, respectively. Qu et al. [21] proposed a fusion rule used for calculating the wavelet transformation modulus maxima of input images with different bandwidths and levels. The approach aims to effectively combine the information from multiple image modalities. Achim et al.[22] developed a fusion rule based on fractional lower-order moments for magnetic resonance (MR) and Computed Tomography (CT) images in the multi-scale medical image analysis literature. Considering the similarities and disparities between the different modalities, Yu et al[23] divided the image into the common component and innovation component and represented the source image with the common and innovation sparse coefficients by joint sparse representation. Wang et al. [24] designed a gate mechanism to dynamically fuse visual features. By utilizing visual features as supplementary information, the proposed method effectively mitigates the influence of noise in voice signals. However, this strategy assumes that the visual data is clean. If the visual data is also corrupted by noise, this solution may be sub-optimal. Sadeghi et al. [25] adopted a latent sequential variable with Markovian dependencies to decide whether visual data is beneficial to audio-visual speech enhancement. Bai et al.[18] proposed a soft-association mechanism to handle inferior image conditions. In practice, the self-attention in Transformer can be used to adaptively decide where and what information should be taken from the image. Cheng et al. [16] adopted both soft attention and hard attention to mitigate noise issue, which is caused by rolling-shutter effect in lidar-stereo fusion task. Specifically, a fusion block is utilized to aggregate supplement information from stereo cameras and generates two disparity maps. Then, by comparing disparity map and lidar point, the noisy points in raw data are filtered out.
@@ 2.1.2 Joint variation based fusion
Joint optimization is often used in variation-based noise reduction when fusing multiple visual modalities, e.g., RGB and thermal. In unimodal noise reduction field, taking image modality as an example, classical denoise model based on Total Variation (TV) is equivalent to solving the following optimization problems:
@@
where $u_{m}(\boldsymbol{x})$ represents the input data from modality $m$, $w_{m}$ is the non-negative weight satisfying $\sum_{m=1}^{M}w_{m}=1$. In a nutshell, joint optimization is the solution of a specific optimization problem using data from different modalities. Wang et al. [26] proposed a variational model in the pixel domain and wavelet domain for joint fusion and denoise of noisy multifocus images. Kumar et al. [27] also employed the total variational model to fuse images acquired using multiple sensors and achieves excellent results on two modal images provided in the medical field and aircraft navigation field. Recently, Padmavathi et al. [28] proposed an image fusion algorithm for constructing a fused image using the total variation model. They incorporated an optimized adaptive weighting fusion scheme for medical multimodal data analysis (i.e., Magnetic Resonance images and Positron Emission Tomography images). Nie et al. [29] introduced a total variation-based fusion method for infrared and visible images. The method leverages the total variation model to fuse two modalities, enabling enhanced image quality and information integration. Quan et al. [30] developed a feature fusion method called relative total variation structure analysis (RTVSA) for urban area classification. This approach combines various features derived from hyperspectral imaging (HSI) and LiDAR data, providing a comprehensive analysis for accurate classification. Furthermore, Liu et al. [31] designed a two-stage enhancement (TSE) framework for infrared and visible image fusion. Their framework incorporated an attention mechanism and a feature-linking model (FLM), combining structure adaptive total-variational (SATV) and L1 sparsity terms to extract two-scale detail layers and a base layer. The proposed method demonstrates robustness and effectiveness in improving the quality of fused images.
@@ 2.2 Cross-modal noise reduction
Many multimodal tasks (e.g., multimodal object detection, vision-and-language understanding) highly depends on correctly aligned multimodal training data. However, the real-world multimodal pairs tend to contain weakly aligned or even unaligned samples [32]. For example, in RGB/thermal multimodal object detection, multimodal inputs are often weakly aligned, i.e., the locations of the same objects may be shifted across different modalities [33]. In social media, the text descriptions are frequently irrelevant with the image content, i.e., unpaired modalities. In this section, we consider weakly aligned or unaligned multimodal samples as cross-modal noise. Compared to modal-specific noise, cross-modal noise lies in a higher-level semantic space. Current cross-modal noise reduction methods can be roughly split into rules-based filtering, model-based rectifying and noise robustness regularization.
@@
Noise robust regularization is another perspective to mitigate the influence of cross-modal noise. To explicitly stabilize and harmonize multimodal pretraining and mitigate the influence of latent noise, NLIP [41] employs noise-adaptive regularization to avoid overfitting to noisy image-text pairs, adjusting alignment labels based on estimated noisy probabilities. Additionally, NLIP utilizes a concept-conditioned cross-modal decoder to generate synthetic captions to impute the missing object information. To mitigate the influence of cross-modal noise in vision-language task, Li et al[42] proposed OSCAR that detects object tags in images and uses them as anchor points for alignment. Recent work provides both theoretical and empirical analysis to shed light on the influence of cross-modal noise in multimodal contrastive learning [43]. Furthermore, the research introduces a novel MMCL loss designed to handle unpaired multimodal samples, demonstrating enhanced robustness in empirical experiments.
@@ 3 Incomplete multimodal Learning
The collected multimodal data in real applications are often incomplete with partial missing modalities for some samples due to unexpected factors, such as equipment damage, data transmission and storage loss. For a concrete example, in user-oriented recommendation system, the browsing behavior history and credit score information may not always be available for some users [48]. Similarly, although combining data from multiple modalities e.g., magnetic resonance imaging (MRI) scans, positron emission tomography (PET) scans, and cerebrospinal fluid (CSF) information, can obtain a more accurate diagnosis for Alzheimer’s Disease [49, 50], owing to high measurement cost for PET scans and uncomfortable invasive tests for CSF, some patients may refuse to conduct these examinations. Thus, incomplete multimodal data commonly exist in Alzheimer’s Disease diagnosis [51]. Typically, conventional multimodal learning models assume the completeness of multimodal data and thus cannot be applied to partial missing modalities directly. For this issue, incomplete multimodal learning, aiming at exploring information of incomplete multimodal data with partial missing modalities, appears and has received increasing research interests in recent years [52]. In this section, we mainly focus on the current progress of researches on incomplete multimodal learning. From the view of whether to impute the missing data, we categorize existing methods into two groups including imputation-based and imputation-free incomplete multimodal learning, where the imputation based methods are further divided into two groups as shown in Fig. 2 including instance and modality level imputation.
@@ 3.1 Imputation based incomplete multimodal learning
For the incomplete multimodal data, a natural approach to address the missing modality learning issue is to directly impute the missing modalities or related missing information in the kernels or graphs constructed from the data as shown in Fig. 2. Then the conventional multimodal learning algorithm can be applied to downstream tasks. In this branch, the existing imputation-based works can be further divided into two categories, including model-agnostic imputation and model-specific imputation. The model-specific imputation methods generally design a imputation model or network for missing data recovery while the imputation model is not necessary for the model-agnostic imputation methods which usually utilizes heuristic filling strategy for the missing modalities.
@@ 3.1.1 Model-agnostic imputation
For the missing modalities, the simplest imputation schemes are zero imputation and mean value imputation, i.e., filling missing values with zero or imputation with the averaged value in the corresponding modality, which is widely used as the baseline methods in most studies [53, 54]. However, previous work [53] shows that the naive imputation may lead to unpromising performance. To address this issue, a weighted matrix factorization model is proposed to assign small weights for the imputed modalities, mitigating the negative influence on optimization [53]. Furthermore, for an online multimodal clustering case, a dynamic weighting and imputation mechanism is designed according to the valid modalities and task at hand[55]. To implement the graph-based approach on incomplete multimodal data, Wang et al[56] utilize similarity values computed from valid modalities to impute the missing elements of graph. In [57], the missing elements in the incomplete kernels are pre-imputed by the average of all columns. In [58], the authors also utilize the average instance of the corresponding modality to impute the missing instances and then assign a small weight to those imputed instances during the graph construction phase.
@@ 3.1.2 Learning based imputation
For the learning based imputation, according to the filling values, we can further divide it into two categories: kernel/graph-based imputation and raw data imputation.
Kernel/graph-based imputation: Kernel and graph learning based approaches are two main lines of incomplete multimodal learning, which transform the learning phase from the raw data to the kernel space or graph. For the kernel and graph constructed from each modality, each element represents the relationship in kernel space or affinity/similarity of instance pairs. The partial missing instances in each modality result in incomplete kernels and graphs. A line of works has been proposed to address the incomplete kernel/graph-based multimodal learning issue. Here we highlight a few seminal works. David Williams et al[59] proposed a Gaussian mixture model based method that can analytically compute the missing Gaussian kernel elements associated with the missing instances. Trivedi et al[60] proposed a kernelized canonical correlation analysis (KCCA) based imputation method, which minimizes a Laplacian kernel regularization to obtain the missing kernel elements. Although effective, it is only applicable to the two-modal data with one complete modality. Lately, collective kernel learning (CKL) [48] is proposed to mutually complete the two incomplete kernels associated with the two incomplete modalities. For the data with more than two modalities, such as the case with $k$ modalities ($k\geq 3$), CKL needs to recover each kernel matrix ${K_{i+1}}$ of the $\left({i+1}\right)$th modality according to kernel ${K_{i}}$ in a cycle iteration scheme. Sahely et al[61] proposed a multimodal kernel completion method based on the sparse reconstruction technique. The method can complete the missing rows and columns of all kernels jointly in a within-kernel preservation and between-kernel exploration framework. Yang et al[62] reduced incomplete multi-modal imputation into semi-supervised learning problem. They further proposed SLIM (Semi-supervised Learning with Incomplete Modality ) and its kernel vision SLIM-K based on matrix completion for solving the modal incompleteness. Recently, Liu et al. proposed several flexible kernel methods to recover the missing kernel elements for incomplete multimodal clustering tasks [63, 64, 65, 66]. For instance, MKKM-IK [63] unifies the kernel completion and consensus partition matrix learning in a unified framework:
@@
By aligning all incomplete kernels to a consensus kernel derived from the consensus orthogonal representation, these methods can jointly obtain the missing elements associated with missing instances and the consensus partition representation shared by all modalities for incomplete multimodal data clustering. Besides of kernel-based methods, an adaptive graph completion method is developed to recover the missing similarities of all graphs with respect to all modalities [67]. The method mainly employs the sparse-representation-based data reconstruction technique to borrow the graph information of other modalities in recovering the missing elements of each modality.
@@
Raw data imputation: Some methods recover the missing modalities at raw feature level. For example, VIGAN (View Imputation via Generative Adversarial Network) [68] is a pioneering work based on generative adversarial network (GAN) and Autoencoder for missing modality recovery. VIGAN employs GAN to initialize the missing modality according to another modality of the same sample and then uses denoise Autoencoder to refine the recovered missing modalities. A limitation of VIGAN is that it is not applicable to data with more than two modalities. Subsequently, partial multimodal clustering via consistent GAN (PMVC_CGAN) [69] and adversarial incomplete multimodal clustering (AIMC) [70] are proposed for missing modality imputation and incomplete multimodal clustering. Different from VIGAN which uses one available data to generate the corresponding missing modality, PMVC_CGAN and AIMC seek to learn the missing modalities from the common representation encoded by the un-missing modalities of the same sample. In [71], a two-stage incomplete deep multimodal network is proposed for breast cancer prediction. Similar to the network structures of AIMC and PMVC_CGAN, GAN is introduced to enhance the quality of missing view imputation. Besides of GAN-based methods, CRA (cascaded residual autoencoder) stacks all modalities into one modal and treats the missing modalities imputation as conventional missing feature completion by optimizing a single-modal cascaded residual autoencoder network [72]. Liu et al[73] proposed another representative autoencoer based imputation framework termed RecFormer for clustering task. Specifically, RecFormer employs a two-stage autoencoder network with the self-attention structure to synchronously extract high-level semantic representations and recover the missing data
@@
To obtain a more reliable modality imputation performance, Tang et al[74] proposed to impute the modality as the average of corresponding semantic neighbors. Lin et al[75] established a dual prediction model that can predict the latent representation of the missing modality according to that of the observed modality and then used the decoder to obtain the recovered missing modality according to the predicted latent representation. matrix factorization is also applied for raw modality imputation. For example, UEAF (Unified Embedding Alignment Framework) [76] and IMCRV (Incomplete multimodal Clustering with Reconstructed Views) [77] provide two matrix factorization based missing modality recovery models, which can reversely reconstruct the data of missing modalities from the common representation. In [78], low-rank representation and tensor learning are introduced for joint missing modality recovery and graph completion. Explicitly taking the reliability of imputation into account, Xie et al[79] introduced the Uncertainty-induced Incomplete Multi-View Data Classification (UIMC) model to address challenges in classifying incomplete multi-view data with missing perspectives. UIMC explores and exploits uncertainty by imputing missing modalities as distributions rather than deterministic values, allowing for more reliable predictions. The model employs an evidential multimodal learning framework to weigh imputed data, reducing the impact of low-quality imputations.